# Charged rotating black holes in scalar multipolar universes

### v1

@@ 4 Conclusions
In our work we further generalized this class of exact solutions of the Einstein-scalar system to include rotation and various charges. More specifically, in section $2$ we presented our solution generating technique that allowed us to combine any stationary exact solution of the Einstein-Maxwell system with axial symmetry with the backreaction of a minimally coupled scalar field. Our technique is based on the symmetries of the dimensionally reduced Lagrangeans down to three dimensions along the timelike direction and it relies on the properties of the Weyl-Papapetrou ansatz for the metric.
As an example we derived the general Kerr-Newman-NUT solution in the scalar multipolar universes. Setting all the charges to zero and for a specific profile of the scalar field we showed how one can recover the rotating version of the Cardoso and Natário solution [6]. One should note that our solution can be further generalized by introducing an acceleration parameter in the original metric [10], [11]. We chose not to include this parameter to keep the final solution listed here in a simpler form.
@@
Another straightforward generalization of our results is to obtain the corresponding solutions in five dimensions, such as the Kerr-Myers solution with scalar fields, or to study the effect of the scalar fields in balancing the static black rings in five dimensions.
@@
Note added After the completion of this work we learned about the paper [24] that presents similar rotating solutions in four and five dimensions for specific scalar profiles.

### v4

@@ 4 Conclusions
In our work we further generalized this class of exact solutions of the Einstein-scalar system to include acceleration, rotation and various other charges. More specifically, in section $2$ we presented our solution generating technique that allowed us to combine any stationary exact solution of the Einstein-Maxwell system with axial symmetry with the backreaction of a minimally coupled scalar field. Our technique is based on the symmetries of the dimensionally reduced Lagrangeans down to three dimensions along the timelike direction and it relies on the properties of the Weyl-Papapetrou ansatz for the metric. Our solution generating technique can be regarded as an alternative way of deriving a result previously known in the literature [9]. However, we do believe that our technique can be easily extended to work in context of general sigma-models, as arising in the dimensional compactification of string theories.
As an example we derived the general accelerating Type D Kerr-Newman-NUT solution in the scalar multipolar universes. In absence of the multipolar scalar field this solution was recently derived in [16] (see also [17], [18]). Setting all the charges to zero and for a specific profile of the scalar field we showed how one can recover the accelerating and rotating version of the Cardoso and Natário solution [6]. If the acceleration parameter is set to zero, $\alpha=0$ one recovers a rotating solution previously derived in [19].
@@
Another straightforward generalization of our results is to obtain the corresponding solutions in five dimensions, such as the Kerr-Myers solution with scalar fields, or to study the effect of the scalar fields in balancing the static black rings in five dimensions. One first step in this direction was taken in [11], where the five dimensional Myers-Perry black hole with a scalar field was first constructed.
Note also that our solution generating technique can be easily extended in presence of other minimally coupled scalar fields (or more generally to a scalar sigma model in four dimensions).